A piece of metal has the shape of a right triangle. The hypotenuse is 3 ft shorter than four times the shortest side and the longer side is 10 ft longer than twice the shorter side. What is the length of the hypotenuse of the​ triangle?

Answer by Boreal(15235)

shorter side=x
longer side is 2x+10
hypotenuse is 4x-3
square them all
x^2+(2x+10)^=(4x-3)^2. Note, binomial expansion (a+b)^2=a^2+2 ab (the internal products are doubled)+b^2
x^4x^2+40x+100=16x^2-24x+9
5x^2+40x+100=16x^2-24x+9
0=11x^2-64x-91
x=(1/22)(64+/- sqrt (4096+4004); sqrt 8100=90
x=(1/22)(64+/-90), use only the + root.
x=(1/22)154=7
sides are 7, 24, with hypotenuse 25, a well known right triangle.